please help me solve this equation: 2 planes which are 1,220 miles apart fly towards eachother, their speeds differ by 60 mph. If they each pass eachother in 2 hours, what is the speed of each plane?
Found 2 solutions by scott8148, mananth:
Answer by scott8148(6628)

their combined speed is such that they close the 1220 gap in 2 hr
___ one is 60 mph faster than the other

x + (x + 60) = 1220 / 2

Answer by mananth(16946)

plane A speed = x mph
plane B speed = x+60 mph
since they are travelling towards each other the effective speed will be 2x+60 mph
...
distance apart 1220 miles
..
time they cross = 2 hours
...
time = distance / speed
1220/2x+60 =2
1220 =2*(2x+60)
1220=4x+120
1220-120=4x
1100=4x
x=1100/4
x=275 mph speed of plane A
Plane B speed = 275+60 = 335 mph